A housing subdivisions has 200 homes. Their growth patter follows the equation y=50x +200, where x is the number of years and y is the number of homes. How many years will it be before reach 650 homes ?
Answer by ankor@dixie-net.com(22740)

A housing subdivisions has 200 homes.
Their growth patter follows the equation y=50x +200, where x is the number of years and y is the number of homes.
How many years will it be before reach 650 homes ?
:
That's y = 650
50x + 200 = 650
50x = 650 - 200
50x = 450
x = 450/50
x = 9 yrs